# Complements of caustics of the real $J_{10}$ singularities

### v3

@@ 6.2. Proof of chirality statements of Theorem 15
If a virtual component of class $\Phi_{1}$ or $\Phi_{3}$ consists of virtual functions with ten real critical points and is achiral, then its $D$-graph has an automorphism, whose restriction to the set of vertices corresponding to minimum points ((or to the set of vertices corresponding to maximum points)) is an involution with at most one fixed point.
Proof. Let $f(x,y)$ be a generic Morse polynomial in an achiral isotopy class with ten real critical points. There are two one-to-one correspondences between its critical points and the critical points of the polynomial f(x,y)f(x,-y). One correspondence preserves the critical values of the critical points, and the other is obtained by continuing over the path connecting the two polynomials in the isotopy class. The composition of these one-to-one correspondences of critical points (and hence of the corresponding vertices of the associated D-graphs) extends to an automorphism of the D-graph of $f$. This automorphism maps the vertex corresponding to the minimum point with the $i$th value from above of coordinate $y$ to the vertex corresponding to the point with the $i$th value from below. Similarly, it permutes the points of maxima. In particular, it defines an involutive permutation of each of these two sets of points with at most one fixed point in each. $\Box$
This proposition implies the chirality of all virtual components with ten real critical points, except for those whose achirality follows from formulas (11) and (19).
It remains to prove the chirality of the component of class $\Phi_{1}$ with eight real critical points, none of which are maxima, see Fig. 18 (right). (The criterion of Theorem 21 does not work for it, since the number of minima is not 2 or 3).
Definition 13.
A level section on a virtual component of the formal graph is a natural-valued function on the set of its vertices, such that
a) for any edge not of type $s2$ this function takes equal values at its endpoints,
b) for any edge corresponding to the $s2$ surgery at which the $i$th and $(i+1)$th real critical values meet and overtake each other, the values at its endpoints are either equal to each other and not equal to $i$ or $i+1$, or one of these values is equal to $i$ and the other to $i+1$.
Theorem 25.
Suppose that $\varphi$ is an arbitrary vertex of an achiral virtual component with an even positive number of minima and the $i$th ((in the ascending order of critical values)) critical point of the virtual function $\varphi$ is a minimum. Then there does not exist a level section of this component, whose value at $\varphi$ is $i$.
Proof. Suppose that such a level section exists. Let $f(x,y)$ be a generic Morse polynomial associated with $\varphi$. Consider again a generic path in the space of Morse functions of class $\Phi_{1}$ or $\Phi_{3}$ connecting the function $f$ with its mirror image f(x,y)f(x,-y), and the path in our virtual component starting from $\varphi$ and consisting of edges corresponding to virtual surgeries describing all real surgeries along the previous path. The latter path is then a cycle. The continuation of the level section’s value along this cycle leads to a number different from the number at its starting point: a contradiction. $\Box$
On the other hand, such a level section for the investigated virtual component was constructed by our program. This finishes the proof of Theorem 15. $\Box$
@@ 7. Proof of Theorem 2
Recall that the canonical Dynkin graphs of the real singularities $D_{4}^{+}$, $D_{6}^{+}$, and $D_{8}^{+}$ are given in Fig. 25.
Figure 25. Real Dynkin graph $D_{8}^{+}$
The canonical Dynkin graphs of the real singularity $D_{2k}^{-}$ is just the standard $D_{2k}$ graph. The canonical Dynkin graphs of the other real simple singularities, $A_{k}$, $D_{2k+1}$, $E_{6}$, $E_{7}$, and $E_{8}$, are also the same as in the classical “complex” theory.
All singularities that appear as perturbations of a $J_{10}$ singularity are simple (or equal to $J_{10}$).
@@
a)) all edges connecting vertices of different subsets are directed from vertices of one subset to vertices of the other, and
b)) the edges of our D-graph, both of whose vertices belong to one of these subsets, form the canonical Dynkin graph of the corresponding real simple singularity $\Xi$ or $\tilde{\Xi}$.
Proof. A proof of the “only if” part repeats the proof of Proposition 26 in [15]. The “if” part essentially follows from the same considerations with the Lyashko–Looijenga covering. (It was not formulated in [15] only because the realizations of all necessary splittings were explicitly presented there). $\Box$
Theorem 2 is a direct corollary of this proposition and the lists of D-graphs given in Theorems 13 and 14. $\Box$

### v5

@@ 6.2. Proof of chirality statements of Theorem 16
If a virtual component of type $\Phi_{1}$ or $\Phi_{3}$ consists of virtual functions with ten real critical points and is achiral, then its $D$-graph has an automorphism, whose restriction to the set of vertices corresponding to minimum points ((or to the set of vertices corresponding to maximum points)) is an involution with at most one fixed point.
Proof. Let $f(x,y)$ be a generic Morse polynomial in an achiral isotopy class with ten real critical points. There are two one-to-one correspondences between its critical points and the critical points of the polynomial f(x,y)f(x,-y). One correspondence preserves the critical values of the critical points, and the other is obtained by tracing the critical points along the path connecting the two polynomials in the isotopy class. The composition of these correspondences of critical points (and, consequently, of the corresponding vertices of the associated D-graphs) extends to an automorphism of the D-graph of $f$. This automorphism maps the vertex corresponding to $i$-th from the right minimum point to the vertex corresponding to the $i$-th from the left minimum point. Similarly, it permutes the maximum points. In particular, it defines an involutive permutation of each of these two sets of points with at most one fixed point in each. $\Box$
This proposition implies the chirality of all virtual components with ten real critical points, listed in Theorems 14 and 15, except for those whose achirality follows from formulas (14) and (23), see Figs. 4 and 10.
It remains to prove the chirality of the isotopy class of type $\Phi_{1}$ with eight real critical points, none of which are maxima, see Fig. 18 (left). (The criterion of Theorem 24 does not work for this class because the number of minima is not 2 or 3).
An isotopy invariant that separates each polynomial $f(x,y)$ of this class from its mirror image f(x,y)f(x,-y) is as follows. There is a single polynomial equation $x=ay^{3}+by^{2}+cy+d$ whose graph in ${\mathbb{R}}^{2}$ contains the four minimum points of $f$. The coefficient $a$ of this polynomial is never zero: otherwise, the restriction of the function $f$ to this curve would be a degree six polynomial in $y$ with four minimum points. Therefore, the sign of this coefficient is an invariant of the isotopy class. These coefficients defined by the functions $f(x,y)$ and f(x,y)f(x,-y) are opposite. $\Box$
This concludes the proof of Theorem 16. $\Box$
@@ 7. Proof of Theorem 2
Recall that the canonical Coxeter-Dynkin graphs of the real singularity classes $D_{4}^{+}$, $D_{6}^{+}$, and $D_{8}^{+}$ are the graphs given in Fig. 27.
Figure 27. Real Coxeter-Dynkin graphs $D_{4}^{+}$, $D_{6}^{+}$, and $D_{8}^{+}$
The canonical Coxeter-Dynkin graph of the real singularity $D_{2k}^{-}$ is just the standard $D_{2k}$ graph. The canonical Coxeter-Dynkin graphs of the other real simple singularities, $A_{k}$, $D_{2k+1}$, $E_{6}$, $E_{7}$, and $E_{8}$, are also the same as in the usual “complex” theory.
All critical points of small perturbations of the $J_{10}$ singularities are simple or belong to the $J_{10}$ class itself.
@@
The following two conditions are equivalent:
  1. (1)
    there exists a polynomial of type $\Phi_{1}$ ((respectively, $\Phi_{3})$ having two real critical points of classes $\Xi$ and $\tilde{\Xi}$ with $\mu(\Xi)+\mu(\tilde{\Xi})=10$;
  2. (2)
@@
    1. (a)
      all edges of the D-graph connecting vertices of different subsets are directed from vertices of the first subset to vertices of the second, and
    2. (b)
      the edges, whose vertices both belong to one of these subsets, form the canonical Coxeter-Dynkin graph of the corresponding real simple singularity $\Xi$ or $\tilde{\Xi}$.
[…]
The condition 1) of Proposition 29 is equivalent to the condition $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{1}$ or $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{3})$ from Theorem 2 because the spaces $\Phi_{1}$ and $\Phi_{3}$ are versal deformations of $J_{10}^{1}$ and $J_{10}^{3}$ singularities, and the group (7) provides the functions with desired critical points arbitrarily close to the origin. Therefore, Theorem 2 follows immediately from Proposition 29 and the lists of D-graphs given in Theorems 14 and 15. $\Box$
8. A problem
To explicitly present the polynomials representing all isotopy classes with ten real critical points. This has only been done above for classes whose $D$-graphs are shown in Figs. 3, 4, 6 (left), 10, 11, 12, and 15 (right).